A stretch limo was about to enter an intersection when it hit its breaks to avoid striking a vehicle already in the intersection. The speed limit is 17.89m/s. The limo traveled a distance ce of17.72 m before colliding into the other vehicle and coming to a stop. It took 4.16 s from when the limos break lights came on until the wreck. What speed was the limo travelling at when it hit its breaks?
1
Expert's answer
2018-11-05T15:53:09-0500

We have:

d=(at^2)/2

The deceleration is

a=2d/t^2 .

The speed was the limo travelling at when it hit its breaks:

v=at=2d/t^2 t=2d/t.

v=2 17.72/(4.16 )=8.52 m/s.